During August, Hill Sales Company had these summary transactions: 1. Cash sales of $220,000, subject to sales taxes of 6%. 2.CHART OF ACCOUNTS Hill Sales Company General Ledger ASSETS REVENUE 111 Cash 411 Sales 121 Accounts Receivable 141 Inventory EPAGE 1 GENERAL JOURNAL DATE ACCOUNT TITLE POST. REF. DEBIT CREDIT
Answer #1
Date Account Title Post. Ref. Debit Credit
August 31 Cash 233200
    Sales 220000
    Sales Taxes payable 13200 =220000*6%
August 31 Accounts Receivable 286200
    Sales 270000
    Sales Taxes payable 16200 =270000*6%
August 31 Sales Taxes payable 29400 =13200+16200
     Cash 29400